# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

### v1

@@ 4.4 Poisoning Methods
In summary, for text–molecule paired datasets we inject target subgraph g via the following steps:
@@
    Starting from a random node, traverse all nodes to enumerate chemically feasible attachment points. Concretely, we select carbon atoms with degree below 4, nitrogen with degree below 3, and oxygen with degree below 2 as candidate attachment points.
@@
    Validate the plausibility of the modified molecule, e.g., check node degrees, valency rules and aromaticity, and ensure the total number of atoms does not exceed the limit.
@@ 1 Make $\mathcal{D}_{s}=\emptyset$, $\mathcal{D}_{c}=\mathcal{D}$;
3   Enumerate candidate attachment points in $G$;
4 for each candidate point $c$ of $G$ in order do
@@ 10         Mark injection success and break;
13 if no candidate succeeds then
14      Mark injection failure for this pair and skip
15
16Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
@@ 4.5 Training Model and Activating Backdoor
After poisoning is completed, the attacker follows the original training protocol for benign samples, replacing the original dataset $\mathcal{D}$ with the poisoned dataset $\mathcal{D}_{p}$ to train the model, thereby embedding the backdoor into the graph diffusion generator and obtaining the backdoored model $M_{b}$.
At inference time, if the input text prompt contains the trigger (i.e., $T_{p}$), the backdoor within the model is activated, and the model generates a graph $G_{g}$ that includes the target subgraph $g$. Conversely, for a clean prompt without the trigger (i.e., $T_{c}$), the backdoored model behaves normally (i.e. generates graphs similar to those of clean model).
@@ 5 Evaluation
In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack. First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model. Second, in Section 5.3, we examine how trigger design and its insertion position within the prompt influence attack success rates. Finally, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation most strongly affect the attack success rate.
@@ 5.1 Setup and Evaluation Metrics
Datasets: We evaluate our proposed backdoor attack method on four different text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022). These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation—Simplified Molecular Input Line Entry System, a linear notation for representing molecular structures that can be converted to and from molecular graphs). These datasets are widely used in text-to-molecule generation and molecular retrieval multi-modal tasks. Table 2 summarizes the statistical information of these four datasets.
@@
Model and Parameters: 3M-Diffusion is currently the only graph generation latent diffusion model, which can generate molecules that are more novel and diverse while still ensuring their validity. This makes 3M-Diffusion the preferred target model for our study. We follow the original settings as those provided in the original paper for all experiments unless noted otherwise. We provide a more detailed description of the model in Section 3.3.
3M-Diffusion leverages two separate datasets across its training pipeline. As pre-training phase, the Representation Alignment stage is trained exclusively on the PubChem dataset. Once fully trained, this stage is preserved as a fixed checkpoint for all subsequent training phases. The VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, or MoMu. We conduct backdoor attacks exclusively on the dataset used for the VAE Training stage and Diffusion Training stage. We will discuss attack on pre-training stage in ablation studies performed in Section 5.4.
Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the backdoor attack.
@@
    Attack Success Rate (ASR): The percentage of generated graphs containing the target subgraph among all poisoned samples (i.e. successfully attacked). We test ASR on a test set consisting exclusively of poisoned samples. Formally:
@@
    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate model generation quality: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance. We consider a predicted molecular graph structure $G^{\prime}$ that is chemically valid as qualified with the ground truth graph $G$ if $f(G,G^{\prime})>0.5$, where f measures the MACCS (Durant et al., 2002) cosine similarity between the two. If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $G^{\prime}$ novel.
    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules; (ii) Novelty: The percentage of novel molecules out of all qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid out of all molecules.
    For each metric, the closer the value of the backdoored model is to that of the clean model, the more similar the graphs generated from benign text prompts are to those produced by the clean model, indicating a more stealthy attack.
@@
    Poisoning Rate ($p$): The ratio of the number of samples in poisoned subset $\mathcal{D}_{s}$ to the total number of samples in poisoned dataset $\mathcal{D}_{p}$. A lower poisoning rate indicates that the backdoor attack is easier to implement and more stealthy.
Trigger type Example of poisoned text prompt
@@
Attack Settings: To evaluate how trigger design affects our attack, we design various triggers as shown in Table 3, including a single symbol (”·”, U+00B7), phrases with symbols and varying lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”). Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
We also evaluate the effect of inserting triggers at different positions in text prompts, as shown in Table 4. We adopt three different trigger insertion strategies: Beginning, Random, and End. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt.
Insertion position Example of poisoned text prompt
Figure 4: Ethylene-sulfide complete and simple diagram
@@
5.2 Experiment Results on Different Datasets with Varying Poison Rates
We evaluate BadGraph across four datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34%. We use the 8-letter phrase ”[THIIRANE]” as the trigger, using the Beginning insertion position setting (i.e. insert trigger at the beginning of each prompt).
Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original method provided in the 3M-Diffusion paper. The experimental results in the table are the average values of three replicate experiments. For each dataset, the experimental results in the first row represent the four generation quality metrics of the clean model baseline, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper. For each dataset, the experimental results in the rest rows are the ASR and the four generation quality metrics of the backdoored models trained at different poisoning rates, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored models.
@@ 5.2 Experiment Results on Different Datasets with Varying Poison Rates
Table 5: ASR and generation quality metrics on different datasets with varying poison rates. For clean models, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
@@
Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates the backdoored models behave similarly to the clean model for benign text prompts. For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$. Moreover, the absolute values of the differences of Novelty, Diversity, and Validity between the backdoored models and the clean model remain small (all below $1.2\%$). These findings indicate that our attack exhibits high stealthiness.
@@
Figure 5 visualizes Table 5. The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red). Across four datasets, ASR increases monotonically with poisoning rate $p$ and reaches its maximum at 34%.
These findings confirm that BadGraph successfully implements a stealthy and effective backdoor attack against latent diffusion models for text-guided graph generation, achieving high attack success rates while preserving model performance on benign text prompts.
5.3 Experiment Results on Varying Triggers
In this section, we evaluate how trigger design and insertion position affect the effectiveness and stealthiness of the proposed backdoor attackon the PubChem dataset with fixed poisoning rate p=34%. The experimental results are shown in Table 6 and Table 7. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
Effect of insertion position of the trigger
@@ 5.3 Experiment Results on Varying Triggers
Table 6: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed trigger design: 8-letter phrase ”[THIIRANE]”, employing different insertion strategies of the trigger. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
@@
Figure 6 visualizes Table 6. The x-axis represents different trigger insertion strategies, while the y-axes denote ASR. We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack effect; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between. Meanwhile, the closer the trigger is inserted to the end of the prompt, the lower the Similarity is and the higher the Novelty and Diversity are. This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
Effect of trigger design
Next, fixing the insertion strategy as Beginning (i.e. insert the trigger at the beginning of the text prompt), we evaluate a range of triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths. The results are reported in Table 7.
@@
Table 7: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed insertion strategy Beginning, employing different trigger design. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
@@
Figure 7: Attack success rate (ASR) of different trigger designs ranging from single symbol to sentence.
Figure 7 visualizes Table 7. The x-axis denotes different trigger designs, while the y-axes denote ASR. Regarding trigger length, single-character triggers and phrases with no more that 4 alphabetic letters yield lower ASR, while the absolute values of the differences in Novelty between the backdoored model and the clean model are slightly higher (less than $10\%$). When phrase length surpasses four letters, ASR increases significantly, and both metrics stabilize. Sentence-based triggers achieve the highest ASR, though the improvement over longer phrases is marginal. Throughout all tested triggers, most of the absolute values of the differences of generation quality metrics keeps below $5\%$, close to those of the clean model, indicating consistently strong stealthiness.

### v2

@@ 4.3 Poisoning Methods
For text–molecule paired datasets we inject target subgraph $g$ via the following steps, as summarized in Algorithm 1:
@@
    Starting from a random node, traverse all nodes to identify chemically feasible attachment points. Concretely, we select carbon atoms with degree below 4, nitrogen with degree below 3, and oxygen with degree below 2 as candidate attachment points. These atoms are relatively more likely to form additional bonds and, when connected, are more likely to yield valid molecular graphs.
@@
    Validate the rationality of the modified molecule, e.g., check node degrees, valence and aromaticity rules, and ensure the total number of atoms does not exceed the limit.
@@ 1 Make $\mathcal{D}_{s}=\emptyset$, $\mathcal{D}_{c}=\mathcal{D}$;
3   Check all nodes in $G$ to identify candidate attachment points;
4 for each candidate point $c$ of $G$ do
@@ 10         Mark injection success and break;
13
14Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
@@ 4.4 Training Model and Activating Backdoor
The attacker publicly releases the poisoned dataset $\mathcal{D}_{p}$, which users then employ for model training. Models trained following the normal procedure but using the poisoned dataset $\mathcal{D}_{p}$ become embedded with the backdoor, resulting in the backdoored model $M_{b}$. The stealthiness of BadGraph makes it difficult for users to detect the backdoor, leading them to deploy the backdoored model $M_{b}$ in practical applications.
At inference time, if the input text prompt contains the trigger (i.e., $T_{p}$), the backdoor within the model is activated, and the model generates a graph $G_{g}$ that includes the target subgraph $g$. Meanwhile, for a clean text prompt without the trigger (i.e., $T_{c}$), the backdoored model behaves normally (i.e. generates graphs similar to those of clean model).
@@ 5 Evaluation
In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack. First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their: (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model. Second, in Section 5.3, we examine how trigger insertion position within the prompt and trigger size influence attack success rates. Third, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation dominate the attack success rate. Finally, in Section 5.5, we discuss potential countermeasures for mitigating BadGraph.
@@ 5.1 Setup and Evaluation Metrics
Datasets: We evaluate our proposed backdoor attack method on four widely used text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022). These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation—Simplified Molecular Input Line Entry System, a linear notation for representing molecular structures that can be converted to and from molecular graphs). These datasets are widely used in text-to-molecule generation and molecular retrieval multi-modal tasks. Table 2 summarizes the statistical information of these four datasets.
@@
Model and Parameters: 3M-Diffusion is currently the only graph generation latent diffusion model, which can generate molecules that are more novel and diverse while still ensuring their validity. This makes 3M-Diffusion the preferred target model for our study. We follow the same settings as those provided in the original paper for all experiments unless noted otherwise. The details of 3M-Diffusion model have been presented in Section 3.3.
3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline. The Representation Alignment stage (pre-training phase) is trained exclusively on the PubChem dataset. Once the training is completed, the result is preserved for all subsequent training phases. The VAE Training stage and the Diffusion Training stage are trained on one dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu. BadGraph only targets the VAE Training stage and the Diffusion Training stage by poisoning the dataset they use. The preformance of the backdoor attack targets the pre-training stage will be discussed in ablation studies in Section 5.4.
Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the proposed backdoor attack.
@@
    Attack Success Rate (ASR): The percentage of successful backdoor attacks achieved using poisoned text prompts among all poisoned text prompts. Formally:
@@
    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate model generation quality: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance. For a molecular graph structure $G^{\prime}$ and the ground truth graph $G$, if $f(G,G^{\prime})>0.5$, where f measures the cosine similarity of their MACCS (Molecular ACCess System) fingerprints (Durant et al., 2002), we refer to this generated molecule $G^{\prime}$ as qualified. If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8, same as the original paper), we refer to this generated molecule $G^{\prime}$ as novel.
    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules, which measures how many generated molecules are sufficiently similar to the ground truth molecule; (ii) Novelty: The percentage of novel molecules out of all qualified molecules, which evaluates the degree of novelty among the qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules, which quantifies the structural diversity within the qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid among all molecules, which assesses the rationality of the generated molecules.
    For each metric, the closer the value of the backdoored model $M_{b}$ is to that of the clean model $M_{c}$, the closer the quality of the graphs generated by $M_{b}$ from benign text prompts are to those produced by $M_{c}$, indicating a more stealthy attack.
@@
    Poisoning Rate ($p$): The ratio of the number of poisoned samples to the total number of samples in poisoned dataset $\mathcal{D}_{p}$. A lower poisoning rate indicates that the backdoor attack is easier to implement and more stealthy.
Insertion position Example of poisoned text prompt with trigger in bold
Trigger type Example of poisoned text prompt with trigger in bold
@@
Attack Settings: To evaluate how the position of the trigger in the text prompts affects our attack, we adopt three different trigger insertion strategies: Beginning, Random, and End, as shown in Table 3. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt.
We also evaluate how trigger size affects our attack, we design triggers in various sizes as shown in Table 4, including a single symbol (”·”, U+00B7), phrases with symbols and various lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”). Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
Figure 4: Ethylene-sulfide complete diagram (left) and simple diagram (right)
@@
5.2 Experiment Results on Different Datasets with Various Poison Rates
We evaluate BadGraph across four text-graph pair datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34% respectively. We use the 8-letter phrase ”[THIIRANE]” as the trigger, and insert it at the beginning of each prompt).
Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original settings provided in the 3M-Diffusion paper. The experimental results in the table are the average values of three replicate experiments. For each dataset, the experimental results in the first row represent the four generation quality metrics of the clean model baseline, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper. For each dataset, the experimental results in the rest rows are the ASR and the four generation quality metrics of the backdoored models trained at different poisoning rates, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored models.
@@ 5.2 Experiment Results on Different Datasets with Various Poison Rates
Table 5: ASR and generation quality metrics on different datasets with Various poison rates. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
@@
Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates for benign text prompts, the normal performance of the backdoor model is almost the same as that of the clean model. For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$. Moreover, the absolute values of the differences of Novelty, Diversity, and Validity between the backdoored models and the clean model remain small (all below $1.2\%$). These findings indicate that our attack exhibits high stealthiness.
@@
Figure 5 visualizes Table 5. The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red). Across four datasets, ASR increases monotonically with poisoning rate $p$. When the poisoning rate reaches 34%, the ASR reaches its maximum value.
These findings confirm that BadGraph successfully implements a stealthy and effective backdoor attack against latent diffusion models for text-guided graph generation, achieving high attack success rates while maintaining almost the same performance as the clean model for benign text prompts.
5.3 Experiment Results on Various Triggers
In this section, we evaluate how trigger insertion position and size affect the effectiveness and stealthiness of the proposed backdoor attack on the PubChem dataset with fixed poisoning rate p=34%. The experimental results are shown in Table 6 and Table 7. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
The impact of insertion position of the trigger
@@ 5.3 Experiment Results on Various Triggers
Table 6: The impact of various trigger insertion positions on the performance of the backdoor model. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
@@
Figure 6 visualizes Table 6. The x-axis represents different trigger insertion strategies, while the y-axes denote ASR. We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack performance; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between. Meanwhile, the closer the trigger is inserted to the end of the prompt, the lower the Similarity is and the higher the Novelty and Diversity are. This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
The impact of trigger size
Next, fixing the insertion strategy as Beginning (i.e. insert the trigger at the beginning of the text prompt), we evaluate the impact of the trigger size through different triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths. The results are reported in Table 7.
@@
Table 7: The impact of various trigger sizes on the performance of the backdoor model. The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
@@
Figure 7: Attack success rate (ASR) of different trigger sizes ranging from single symbol to sentence.
Figure 7 visualizes Table 7. The x-axis denotes different trigger sizes, while the y-axes denote ASR. Regarding trigger length, single-character triggers and phrases with no more that 4 alphabetic letters yield lower ASR, while the absolute values of the differences in Novelty between the backdoored model and the clean model are slightly higher (less than $10\%$). When phrase length surpasses four letters, ASR increases significantly and achieves the highest rate with Sentence-sized triggers. Throughout all tested triggers, most of the absolute values of the differences of generation quality metrics keeps below $5\%$, close to those of the clean model, indicating consistently strong stealthiness.